# supplementary

### v1

@@ Abstract
Learning robust and generalizable world models is crucial for enabling efficient and scalable robotic control in real-world environments. In this work, we introduce a novel framework for learning world models that accurately capture complex, partially observable, and stochastic dynamics. The proposed method employs a dual-autoregressive mechanism and self-supervised training to achieve reliable long-horizon predictions without relying on domain-specific inductive biases, ensuring adaptability across diverse robotic tasks. We further propose a policy optimization framework that leverages world models for efficient training in imagined environments and seamless deployment in real-world systems. Through extensive experiments, our approach consistently outperforms state-of-the-art methods, demonstrating superior autoregressive prediction accuracy, robustness to noise, and generalization across manipulation and locomotion tasks. Notably, policies trained with our method are successfully deployed on ANYmal D hardware in a zero-shot transfer, achieving robust performance with minimal sim-to-real performance loss. This work advances model-based reinforcement learning by addressing the challenges of long-horizon prediction, error accumulation, and sim-to-real transfer. By providing a scalable and robust framework, the introduced methods pave the way for adaptive and efficient robotic systems in real-world applications.
I Introduction
Robotic systems have achieved remarkable advancements in recent years, driven by progress in reinforcement learning (RL) [1, 2] and control theory [3, 4]. High-fidelity simulators have played a pivotal role in these developments, enabling safe and efficient policy training [4, 5, 6]. However, the transfer of policies from simulation to real-world environments remains a significant challenge due to inevitable inaccuracies in modeling real-world dynamics [7, 8, 9]. These inaccuracies lead to performance degradation when deploying learned policies on physical hardware, a problem commonly referred to as the sim-to-real gap.
@@ I Introduction
A prevalent limitation in many approaches is the lack of adaptation and learning once the policy is deployed on the real system [10, 11, 12]. This results in underutilization of the valuable data generated during real-world interactions. Robotic systems operating in dynamic and uncertain environments require the ability to continually adapt their behavior to new conditions. The inability to exploit real-world experience for further learning restricts the system’s robustness and limits its ability to handle evolving scenarios effectively. Truly intelligent robotic systems should operate efficiently and reliably using limited data, adapting to real-world conditions in a scalable manner [13, 14]. While model-free RL algorithms such as Proximal Policy Optimization (PPO) [2] and Soft Actor-Critic (SAC) [1] have demonstrated impressive results in simulation, their high interaction requirements make them impractical for real-world robotics. Sample-efficient methods are therefore essential for leveraging the information in real-world data without extensive environment interactions [15, 16].
A promising solution is the use of predictive models of the environment, commonly referred to as world models[17, 18]. World models simulate environment dynamics to enable planning and policy optimization, often referred to as learning in imagination [19]. These models have shown success across diverse robotic domains, including manipulation [20, 21], navigation [14], and locomotion [13]. However, developing reliable and generalizable world models poses unique challenges due to the complexity of real-world dynamics, including nonlinearities, stochasticity, and partial observability [22, 23].
@@
Existing approaches often incorporate domain-specific inductive biases, such as structured state representations or hand-designed network architectures [24, 25, 26], to improve model fidelity. While effective, these methods are limited in their scalability and adaptability to novel environments or tasks. In contrast, a general framework for learning world models without domain-specific assumptions has the potential to enhance generalization and applicability across a wide range of robotic systems and scenarios.
@@
Our contributions are summarized as follows:
  • We introduce a novel network architecture and training framework that enables the learning of reliable world models capable of long autoregressive rollouts, a critical property for downstream planning and control.
  • We provide a comprehensive evaluation suite spanning diverse robotic tasks to benchmark our method. Comparative experiments with existing world model frameworks demonstrate the effectiveness of our approach.
  • We propose an efficient policy optimization framework that leverages the learned world models for continuous control and generalizes effectively to real-world scenarios.
@@
By addressing the challenges associated with learning world models, this work contributes toward bridging the gap between data-driven modeling and real-world deployment. The proposed framework enhances the scalability, adaptability, and robustness of robotic systems, paving the way for broader adoption of model-based reinforcement learning in real-world applications.
II Related work
II-A World Models for Robotics
@@
World models have emerged as a cornerstone in robotics for capturing system dynamics and enabling efficient planning and control through simulated trajectories. Early works like PILCO [27] demonstrated exceptional sample efficiency by leveraging Gaussian processes, but their applicability to high-dimensional, nonlinear systems has been limited by scalability constraints. The advent of deep neural networks expanded the horizons of dynamics modeling, allowing complex robotic systems to be effectively modeled and controlled. For example, neural network dynamics models have been integrated with Model Predictive Control (MPC) to enhance adaptability in real-world settings [13]. A prominent application of world models is in robotic control, where dynamics models are used to describe real-world dynamics for policy optimization [28]. Extensions to vision-based tasks have been realized through visual foresight techniques [21, 29, 20], which learn visual dynamics for planning in high-dimensional sensory spaces. Similar ideas are applied to train RL agents in such world models aiming to fully replicate real environment interactions [17, 30]. These approaches underline the versatility of world models in tasks requiring rich perceptual inputs.
@@
To improve the generalization of black-box neural network-based world models beyond the training distribution, many works incorporate known physics principles or state structures into model design, addressing potential limitations in control performance. Examples include foot-placement dynamics [24], object invariance [25], granular media interactions [31], frequency domain parameterization [26], rigid body dynamics [23], and semi-structured Lagrangian dynamics models [32]. While these methods demonstrate impressive results, they often require strong domain knowledge and carefully crafted inductive biases, which can restrict their scalability and adaptability to diverse robotic applications. Latent-space dynamics models offer an alternative by abstracting the state space into compact representations, enabling efficient long-horizon planning. Deep Planning Network (PlaNet) [18] and its successor Dreamer [33, 14, 34] exemplify this trend, achieving state-of-the-art performance in continuous control and visual navigation tasks. These frameworks have been extended to real-world robotics [22, 35], demonstrating their potential in both simulation and hardware deployment.
II-B Model-Based Reinforcement Learning
@@
Model-Based Reinforcement Learning (MBRL) has emerged as a powerful approach to address the limitations of model-free reinforcement learning, particularly in scenarios where sample efficiency and safety are critical. Unlike model-free methods, which learn policies directly from interactions with the environment, MBRL leverages a learned model of the environment to simulate interactions, enabling more efficient and safer policy learning. One of the pioneering methods in MBRL is Probabilistic Ensembles with Trajectory Sampling (PETS), which uses an ensemble of probabilistic neural networks to model the environment dynamics [15]. This ensemble approach captures the uncertainty in the environment, allowing for more robust decision-making. Building on the idea of latent-space modeling, PlaNet leverages a latent dynamics model to plan directly in a learned latent space [18]. This approach enables long-horizon planning and is particularly effective in environments where direct planning in the raw state space would be computationally prohibitive. Similar ideas were also seen in previous attempts [36, 37]. Dreamer extends the concept by incorporating an actor-critic framework into the latent dynamics model, enabling the simultaneous learning of both the dynamics model and the policy [33, 14, 34]. Variations on the architectural design also see success in improving generation capabilities of such latent dynamics models with autoregressive transformer [38] and the stochastic nature of variational autoencoders [39]. Recent advancements in this area include TD-MPC and TD-MPC2, which integrate model-based learning with MPC to achieve high-performance control in dynamic environments. TD-MPC leverages temporal difference learning to refine the model and improve the accuracy of long-term predictions [40, 41], while TD-MPC2 extends this approach by incorporating more sophisticated planning strategies to handle environments with high variability [42].
@@
Recognizing the strengths of both model-based and model-free methods, several hybrid approaches have been developed to combine the sample efficiency of MBRL with the robustness of model-free reinforcement learning. One notable example is Model-Based Policy Optimization (MBPO), which uses a model-based approach for planning and policy optimization but refines the policy using model-free updates [16]. It emphasizes selectively relying on the learned model when its predictions are accurate, thus mitigating the negative effects of model inaccuracies. Building on similar principles, Model-based Offline Policy Optimization (MOPO) extends the framework to the offline setting, where learning is conducted entirely from previously collected data without further environment interaction [43]. MOPO introduces uncertainty penalization in the learned model’s predictions to ensure conservative policy updates, effectively reducing the risk of overfitting to model errors. Model-Based Meta-Policy-Optimization (MB-MPO) advances MBPO by incorporating meta-learning techniques to improve the generalization of learned policies across different tasks [44]. This method trains a meta-policy that can quickly adapt to new environments by leveraging a shared model, enhancing the applicability of MBPO in multi-task settings. In contrast to using zeroth-order model-free reinforcement learning for policy optimization, first-order gradient-based optimization is used to improve policy learning [45, 46]. These methods scale up the world models to capture intricate details of environment dynamics and directly apply gradient-based optimization to improve policy learning. This allows for more efficient and precise policy updates, particularly in complex, high-dimensional environments, where accurate gradient information is crucial for performance.
III Approach
III-A Reinforcement Learning and World Models
@@
We formulate the problem by modeling the environment as a Partially Observable Markov Decision Process (POMDP) [47], defined by the tuple $\left({\mathcal{S}},{\mathcal{A}},{\mathcal{O}},T,R,O,\gamma\right)$, where ${\mathcal{S}}$, ${\mathcal{A}}$, and ${\mathcal{O}}$ denote the state, action, and observation spaces, respectively. The transition kernel $T:{\mathcal{S}}\times{\mathcal{A}}\to{\mathcal{S}}$ captures the environment dynamics $p\left(s_{t+1}\mid s_{t},a_{t}\right)$, while the reward function $R:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to{\mathbb{R}}$ maps transitions to scalar rewards. Observations $o_{t}\in{\mathcal{O}}$ are emitted according to probabilities $p\left(o_{t}\mid s_{t}\right)$, governed by the observation function $O:\mathcal{S}\to\mathcal{O}$. The agent seeks to learn a policy $\pi_{\theta}:{\mathcal{O}}\to{\mathcal{A}}$ that maximizes the expected discounted return $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t\geq 0}\gamma^{t}r_{t}\right]$, where $r_{t}$ is the reward at time $t$ and $\gamma\in\left[0,1\right]$ is the discount factor.
@@
World models [17] approximate the environment dynamics by learning $p\left(s_{t+1}\mid s_{t},a_{t}\right)$. These models facilitate policy optimization by enabling simulated environment interactions in imagination [19]. Training typically involves three iterative steps: (1) collect data from real environment interactions; (2) train the world model using the collected data; and (3) optimize the policy within the simulated environment produced by the world model.
@@
III-B Self-supervised Autoregressive Training
@@
Training data is constructed by sliding a window of size $M+N$ over collected trajectories, providing sufficient historical context for prediction targets. To improve gradient propagation through autoregressive predictions, we apply reparameterization tricks to enable effective end-to-end optimization. By incorporating historical observations, RWM captures unobservable dynamics, addressing the challenges of partially observable and potentially discontinuous environments. The autoregressive training mitigates error accumulation, a common issue in long-horizon predictions, and eliminates the need for handcrafted representations or domain-specific inductive biases, enhancing generalization across diverse tasks. This process is illustrated in Fig. 0(a), in contrast to the teacher-forcing pipeline in Fig. 0(b), which is commonly adopted to train many popular architectures [33, 48]. Specifically, teacher-forcing can be viewed as a special case of autoregressive training with forecast horizon $N=1$, which boosts training with higher parallelization.
@@
While the proposed autoregressive training framework can be applied to any network architecture, RWM utilizes a GRU-based architecture for its ability to maintain long-term historical context while operating on low-dimensional inputs. The network predicts the mean and standard deviation of a Gaussian distribution describing the next observation. Our framework introduces a dual-autoregressive mechanism:
  • Inner autoregression updates GRU hidden states autoregressively after each historical step within the context horizon $M$.
  • Outer autoregression feeds predicted observations from the forecast horizon $N$ back into the network.
This architecture, visualized in Fig. 1, ensures robustness to long-term dependencies and transitions, making RWM suitable for complex robotics applications.
Figure 1: Dual-autoregressive mechanism employed in RWM. Inner autoregression updates GRU hidden states after each historical step within the context horizon, while outer autoregression feeds predicted observations from the forecast horizon back into the network. The dashed arrows denote the sequential autoregressive prediction steps, highlighting robustness to long-term dependencies and transitions.
III-C Policy Optimization on Learned World Models
Policy optimization in RWM is conducted using the learned world model, following a framework inspired by Model-Based Policy Optimization (MBPO) [16] and the Dyna algorithm [49]. During imagination, the actions are generated recursively by the policy $\pi_{\theta}$ conditioned on the observations predicted by the world model $p_{\phi}$, which is further conditioned on the previous predictions. The actions at time $t+k$ can thus be written as
@@ III-C Policy Optimization on Learned World Models
where $o^{\prime}_{t+k}$ is drawn autoregressively from the distribution according to Eq. 1. The approach combines model-based imagination with model-free reinforcement learning to achieve efficient and robust policy optimization, as outlined in Algorithm 1.
@@
2:  for learning iterations $=1,2,\dots$ do
@@
4:     Update the world model $p_{\phi}$ with autoregressive training using data sampled from ${\mathcal{D}}$ according to Eq. 2
@@
The replay buffer ${\mathcal{D}}$ aggregates real environment interactions collected by a single agent. The world model $p_{\phi}$ is trained on this data following the autoregressive scheme described in Sec. III-B. Imagination agents are initialized from samples in ${\mathcal{D}}$ and simulate trajectories using the world model for $T$ steps, enabling policy updates through a reinforcement learning algorithm. The training diagram is visualized in Fig. 2.
Figure 2: Model-Based Policy Optimization with learned world models. The framework combines real environment interactions with simulated rollouts for efficient policy optimization. Observation and action pairs from the environment are stored in a replay buffer and used to train the autoregressive world model. Imagination rollouts using the learned model predict future states over a horizon of $T$, providing trajectories for policy updates through reinforcement learning algorithms.
@@
While PPO is known for its strong performance in robotic tasks, training it on learned world models poses unique challenges. Model inaccuracies can be exploited during policy learning, leading to discrepancies between the imagined and true dynamics. This issue is exacerbated by the extended autoregressive rollouts required for PPO, which compound prediction errors. We denote this policy optimization method by MBPO-PPO.
Despite these challenges, RWM demonstrates its robustness by successfully optimizing policies over a hundred autoregressive steps with MBPO-PPO, far exceeding the capabilities of existing frameworks such as MBPO [16], Dreamer [33, 14, 34], or TD-MPC [40, 42]. This result underscores the accuracy and stability of the proposed training method and its ability to synthesize policies deployable on hardware. By using MBPO-PPO, we provide a stringent test of model fidelity, establishing the effectiveness of our framework for real-world robotic tasks.
IV Experiments

### v4

# Robotic World Model: A Neural Network Simulator for Robust Policy Optimization in Robotics

@@ Abstract
Learning robust and generalizable world models is crucial for enabling efficient and scalable robotic control in real-world environments. In this work, we introduce a novel framework for learning world models that accurately capture complex, partially observable, and stochastic dynamics. The proposed method employs a dual-autoregressive mechanism and self-supervised training to achieve reliable long-horizon predictions without relying on domain-specific inductive biases, ensuring adaptability across diverse robotic tasks. We further propose a policy optimization framework that leverages world models for efficient training in imagined environments and seamless deployment in real-world systems. This work advances model-based reinforcement learning by addressing the challenges of long-horizon prediction, error accumulation, and sim-to-real transfer. By providing a scalable and robust framework, the introduced methods pave the way for adaptive and efficient robotic systems in real-world applications.
Figure 1: Autoregressive imagination, ground-truth simulation, and real-world deployment of RWM. For each environment, the top row showcases the RWM autoregressively predicting future trajectories in imagination. The second row visualizes the ground truth evolution in simulation. Specifically for the ANYmal D quadruped and Unitree G1 humanoid, the framework achieves robust policy optimization through MBPO-PPO, enabling zero-shot deployment on hardware.
1 Introduction
@@
Robotic systems have achieved remarkable advancements in recent years, driven by progress in reinforcement learning (RL) [haarnoja2018soft, schulman2017proximal] and control theory [nguyen2019review, todorov2012mujoco]. A prevalent limitation in many approaches is the lack of adaptation and learning once the policy is deployed on the real system [tan2018sim, peng2020learning, li2023learning, li2023versatile]. This results in underutilization of the valuable data generated during real-world interactions. Robotic systems operating in dynamic and uncertain environments require the ability to continually adapt their behavior to new conditions [lee2020learning]. The inability to exploit real-world experience for further learning restricts the system’s robustness and limits its ability to handle evolving scenarios effectively. Truly intelligent robotic systems should operate efficiently and reliably using limited data, adapting to real-world conditions in a scalable manner [nagabandi2018neural, hafner2020mastering]. While model-free RL algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximal] and Soft Actor-Critic (SAC) [haarnoja2018soft] have demonstrated impressive results in simulation, their high interaction requirements make them impractical for real-world robotics. Sample-efficient methods are therefore essential for leveraging the information in real-world data without extensive environment interactions [chua2018deep, janner2019trust].
@@
A promising solution is the use of predictive models of the environment, commonly referred to as world models [ha2018recurrent, hafner2019learning]. World models simulate environment dynamics to enable planning and policy optimization, often referred to as learning in imagination [sutton1991dyna]. These models have shown success across diverse robotic domains, including manipulation [ebert2018visual, finn2016unsupervised], navigation [hafner2020mastering], and locomotion [nagabandi2018neural]. However, developing reliable and generalizable world models poses unique challenges due to the complexity of real-world dynamics, including nonlinearities, stochasticity, and partial observability [wu2023daydreamer, song2024learning]. Existing approaches often incorporate domain-specific inductive biases, such as structured state representations or hand-designed network architectures [yang2020data, sancaktar2022curious, li2024fld], to improve model fidelity. While effective, these methods are limited in their scalability and adaptability to novel environments or tasks. In contrast, a general framework for learning world models without domain-specific assumptions has the potential to enhance generalization and applicability across a wide range of robotic systems and scenarios.
@@
Our contributions are summarized as follows: (i) We introduce a novel network architecture and training framework that enables the learning of reliable world models capable of long autoregressive rollouts, a critical property for downstream planning and control. (ii) We provide a comprehensive evaluation suite spanning diverse robotic tasks to benchmark our method. Comparative experiments with existing world model frameworks demonstrate the effectiveness of our approach. (iii) We propose an efficient policy optimization framework that leverages the learned world models for continuous control and generalizes effectively to real-world scenarios with hardware experiments, including both quadruped and humanoid systems.
@@
By addressing the challenges associated with learning world models, this work contributes toward bridging the gap between data-driven modeling and real-world deployment. The proposed framework enhances the scalability, adaptability, and robustness of robotic systems, paving the way for broader adoption of model-based reinforcement learning in real-world applications. Supplementary videos for this work are available on https://sites.google.com/view/roboticworldmodel.
2 Related work
2.1 World Models for Robotics
@@
World models have emerged as a cornerstone in robotics for capturing system dynamics and enabling efficient planning and control through simulated trajectories. A prominent application of world models is in robotic control, where dynamics models are used to describe real-world dynamics for policy optimization [levine2016end]. Extensions to vision-based tasks have been realized through visual foresight techniques [finn2016unsupervised, finn2017deep, ebert2018visual], which learn visual dynamics for planning in high-dimensional sensory spaces. Similar ideas are applied to train RL agents in such world models aiming to fully replicate real environment interactions [ha2018recurrent, alonso2024diffusion]. These approaches underline the versatility of world models in tasks requiring rich perceptual inputs.
@@
To improve the generalization of black-box neural network-based world models beyond the training distribution, many works incorporate known physics principles or state structures into model design, addressing potential limitations in control performance. Examples include foot-placement dynamics [yang2020data], object invariance [sancaktar2022curious], granular media interactions [choi2023learning], frequency domain parameterization [li2024fld], rigid body dynamics [song2024learning], and semi-structured Lagrangian dynamics models [levy2024learning]. While these methods demonstrate impressive results, they often require strong domain knowledge and carefully crafted inductive biases, which can restrict their scalability and adaptability to diverse robotic applications. Latent-space dynamics models offer an alternative by abstracting the state space into compact representations, enabling efficient long-horizon planning. Deep Planning Network (PlaNet) [hafner2019learning] and its successor Dreamer [hafner2019dream, hafner2020mastering, hafner2023mastering] exemplify this trend, achieving state-of-the-art performance in continuous control and visual navigation tasks. These frameworks have been extended to real-world robotics [wu2023daydreamer, bi2024sample], demonstrating their potential in both simulation and hardware deployment.
2.2 Model-Based Reinforcement Learning
@@
Model-Based Reinforcement Learning (MBRL) has emerged as a powerful approach to address the limitations of model-free reinforcement learning, particularly in scenarios where sample efficiency and safety are critical. Unlike model-free methods, which learn policies directly from interactions with the environment, MBRL leverages a learned model of the environment to simulate interactions, enabling more efficient and safer policy learning. One of the pioneering methods in MBRL is Probabilistic Ensembles with Trajectory Sampling (PETS), which uses an ensemble of probabilistic neural networks to model the environment dynamics [chua2018deep]. Building on the idea of latent-space modeling, PlaNet leverages a latent dynamics model to plan directly in a learned latent space [hafner2019learning]. Dreamer extends the concept by incorporating an actor-critic framework into the latent dynamics model, enabling the simultaneous learning of both the dynamics model and the policy [hafner2019dream, hafner2020mastering, hafner2023mastering]. Variations on the architectural design also see success in improving generation capabilities of such latent dynamics models with autoregressive transformer [micheli2022transformers] and the stochastic nature of variational autoencoders [zhang2024storm]. Recent advancements in this area include TD-MPC and TD-MPC2, which integrate model-based learning with MPC to achieve high-performance control in dynamic environments [hansen2022temporal, feng2023finetuning, hansen2023td].
@@
Recognizing the strengths of both model-based and model-free methods, several hybrid approaches have been developed to combine the sample efficiency of MBRL with the robustness of model-free reinforcement learning. One notable example is Model-Based Policy Optimization (MBPO), which uses a model-based approach for planning and policy optimization but refines the policy using model-free updates [janner2019trust]. It emphasizes selectively relying on the learned model when its predictions are accurate, thus mitigating the negative effects of model inaccuracies. Building on similar principles, Model-based Offline Policy Optimization (MOPO) extends the framework to the offline setting, where learning is conducted entirely from previously collected data without further environment interaction [yu2020mopo]. In contrast to using zeroth-order model-free reinforcement learning for policy optimization, first-order gradient-based optimization is used to improve policy learning [xu2022accelerated, georgiev2024adaptive]. This allows for more efficient and precise policy updates, particularly in complex, high-dimensional environments, where accurate gradient information is crucial for performance. Our framework extends MBPO by integrating it with PPO over extensive autoregressive rollouts, making it particularly effective for complex robotic control tasks.
3 Approach
3.1 Reinforcement Learning and World Models
@@
We formulate the problem by modeling the environment as a Partially Observable Markov Decision Process (POMDP) [sutton2018reinforcement], defined by the tuple $\left({\mathcal{S}},{\mathcal{A}},{\mathcal{O}},T,R,O,\gamma\right)$, where $\mathcal{S}$}, $\mathcal{A}$}, and $\mathcal{O}$} denote the state, action, and observation spaces, respectively. The transition kernel $T:{\mathcal{S}}\times{\mathcal{A}}\to{\mathcal{S}}$ captures the environment dynamics $p\left(s_{t+1}\mid s_{t},a_{t}\right)$, while the reward function $R:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to{\mathbb{R}}$ maps transitions to scalar rewards. Observations $o_{t}\in{\mathcal{O}}$ are emitted according to probabilities $p\left(o_{t}\mid s_{t}\right)$, governed by the observation kernel $O:\mathcal{S}\to\mathcal{O}$. The agent seeks to learn a policy $\pi_{\theta}:{\mathcal{O}}\to{\mathcal{A}}$ that maximizes the expected discounted return $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t\geq 0}\gamma^{t}r_{t}\right]$, where $r_{t}$ is the reward at time $t$ and $\gamma\in\left[0,1\right]$ is the discount factor.
@@
World models [ha2018recurrent] approximate the environment dynamics and facilitate policy optimization by enabling simulated environment interactions in imagination [sutton1991dyna]. Training typically involves three iterative steps: (1) collect data from real environment interactions; (2) train the world model using the collected data; and (3) optimize the policy within the simulated environment produced by the world model.
@@
3.2 Self-supervised Autoregressive Training
@@
Training data is constructed by sliding a window of size $M+N$ over collected trajectories, providing sufficient historical context for prediction targets. To improve gradient propagation through autoregressive predictions, we apply reparameterization tricks to enable effective end-to-end optimization. By incorporating historical observations, RWM captures unobservable dynamics, addressing the challenges of partially observable and potentially discontinuous environments. The autoregressive training mitigates error accumulation, a common issue in long-horizon predictions, and eliminates the need for handcrafted representations or domain-specific inductive biases, enhancing generalization across diverse tasks. This process is illustrated in Fig. 2(a), in contrast to the teacher-forcing pipeline in Fig. 2(b), which is commonly adopted to train many popular architectures [hafner2019dream, chen2021decision]. Specifically, teacher-forcing can be viewed as a special case of autoregressive training with forecast horizon $N=1$, which boosts training with higher parallelization.
@@
While the proposed autoregressive training framework can be applied to any network architecture, RWM utilizes a GRU-based architecture for its ability to maintain long-term historical context while operating on low-dimensional inputs. The network predicts the mean and standard deviation of a Gaussian distribution describing the next observation. Our framework introduces a dual-autoregressive mechanism: (i) Inner autoregression updates GRU hidden states autoregressively after each historical step within the context horizon $M$. (ii) Outer autoregression feeds predicted observations from the forecast horizon $N$ back into the network. This architecture, visualized in Fig. S6, ensures robustness to long-term dependencies and transitions, making RWM suitable for complex robotics applications.
3.3 Policy Optimization on Learned World Models
Policy optimization in RWM is conducted using the learned world model, following a framework inspired by Model-Based Policy Optimization (MBPO) [janner2019trust] and the Dyna algorithm [sutton1990integrated]. During imagination, the actions are generated recursively by the policy $\pi_{\theta}$ conditioned on the observations predicted by the world model $p_{\phi}$, which is further conditioned on the previous predictions. The actions at time $t+k$ can thus be written as
@@ 3.3 Policy Optimization on Learned World Models
where $o^{\prime}_{t+k}$ is drawn autoregressively according to Eq. 1. Rewards are computed from imagined observations and privileged information. The approach combines model-based imagination with model-free RL to achieve efficient and robust policy optimization, as outlined in Algorithm 1.
@@
2:for learning iterations $1,2,\dots$ do
@@
4:  Update $p_{\phi}$ with autoregressive training using data sampled from $\mathcal{D}$} according to Eq. 2
@@
The replay buffer $\mathcal{D}$} aggregates real environment interactions collected by a single agent. The world model $p_{\phi}$ is trained on this data following the autoregressive scheme described in Sec. 3.2. Imagination agents are initialized from samples in $\mathcal{D}$} and simulate trajectories using the world model for $T$ steps, enabling policy updates through a reinforcement learning algorithm. The training diagram is visualized in Fig. S7.
@@
While PPO is known for its strong performance in robotic tasks, training it on learned world models poses unique challenges. Model inaccuracies can be exploited during policy learning, leading to discrepancies between the imagined and true dynamics. This issue is exacerbated by the extended autoregressive rollouts required for PPO, which compound prediction errors. We denote this policy optimization method by MBPO-PPO. Despite these challenges, RWM demonstrates its robustness by successfully optimizing policies over a hundred autoregressive steps with MBPO-PPO, far exceeding the capabilities of existing frameworks such as MBPO [janner2019trust], Dreamer [hafner2019dream, hafner2020mastering, hafner2023mastering], or TD-MPC [hansen2022temporal, hansen2023td]. This result underscores the accuracy and stability of the proposed training method and its ability to synthesize policies deployable on hardware.
4 Experiments